# Adapt On-the-Go: Behavior Modulation for Single-Life Robot Deployment

### v2

@@ A.4 General Experiment Setup Details
where q distance is between 0 and 1 and decays quadratically until a threshold which is the robot’s action limits. Specifically, we follow the reward structure put forth by tunyasuvunakool2020. These terms comprise the overwhelming majority of weight in the final reward. We also include terms for avoiding undesirable behaviors like rocking or swaying that penalize any angular velocity in the root body’s roll, pitch, and yaw. We also slightly penalize energy consumption and torque smoothness. To encourage a walking gait in particular, we added another regularization term to encourage diagonal shoulder and hip joints to be the same at any given time.
A.5 Implementation Details and Hyperparameters
We implemented all methods, including ROAM, RMA, and HLC, on top of the same state-of-the art implementation of SAC from smith2022walk as the base learning approach. For all comparisons, we additionally use a high UTD ratio, dropout, and layernorm, following DroQ [hiraoka2021dropout], and for methods that do online fine-tuning, we use 50/50 sampling following RLPD [ball2023rlpd]. We use default hyperparameter values: a learning rate of $3\times 10^{-4}$, an online batch size of $128$, and a discount factor of $0.99$. The policy and critic networks are MLPs with 2 hidden layers of 256 units each and ReLU activations. For ROAM, we tuned $\beta$ with values 0.01, 0.1, 0.5, 0.9.
@@ A.5 Implementation Details and Hyperparameters
For the real-world experiments, the state space consists of joint positions, joint velocities, torques, forward linear velocity, IMU (roll, pitch, change in roll, change in pitch), and normalized foot forces for a total of $47$ dimensions. We use an Intel T265 camera-based velocity estimator to estimate onboard linear velocity. We use $K_{p}$ and $K_{d}$ gains of $20$ and $1$, respectively, which are used in the position controller. We again use action interpolation, an action range of 35% physical limits, and a $1$ step action history. We also use a second-order Butterworth low-pass filter with a high-cut value of $8$ to smooth the position targets. Finally, to reset the robot, we use the reset policy provided by smith2022walk. We train $4$ prior behavior policies for the real-world experiments, each of which is trained with a frozen knee joint. Specifically, we train a policy with the front right knee joint frozen, the front left knee joint frozen, the rear right knee joint frozen, and the rear left knee joint frozen. $\beta=0.5$ is used in all real-world experiments for ROAM.
@@
For RMA training, we changed the environment dynamics between each episode and trained for a total of $2,000,000$ iterations. The standard architecture and hyperparameter choices from [kumar2021rma] were used along with DroQ [hiraoka2021dropout] as the base algorithm.

### v3

@@ A.3 General Experiment Setup Details
where q distance is between 0 and 1 and decays quadratically until a threshold which is the robot’s action limits. Specifically, we follow the reward structure put forth by (Tunyasuvunakool et al., 2020). These terms comprise the overwhelming majority of weight in the final reward. We also include terms for avoiding undesirable behaviors like rocking or swaying that penalize any angular velocity in the root body’s roll, pitch, and yaw. We also slightly penalize energy consumption and torque smoothness. To encourage a walking gait in particular, we added another regularization term to encourage diagonal shoulder and hip joints to be the same at any given time.
A.4 Implementation Details and Hyperparameters
We implemented all methods, including ROAM, RMA, and HLC, on top of the same state-of-the art implementation of SAC from (Smith et al., 2022) as the base learning approach. For all comparisons, we additionally use a high UTD ratio, dropout, and layernorm, following DroQ (Hiraoka et al., 2021), and for methods that do online fine-tuning, we use 50/50 sampling following RLPD (Ball et al., 2023). We use default hyperparameter values: a learning rate of $3\times 10^{-4}$, an online batch size of $128$, and a discount factor of $0.99$. The policy and critic networks are MLPs with 2 hidden layers of 256 units each and ReLU activations. For ROAM, we tuned $\beta$ with values 0.01, 0.1, 0.5, 0.9.
@@ A.4 Implementation Details and Hyperparameters
For the real-world experiments, the state space consists of joint positions, joint velocities, torques, forward linear velocity, IMU (roll, pitch, change in roll, change in pitch), and normalized foot forces for a total of $47$ dimensions. We use an Intel T265 camera-based velocity estimator to estimate onboard linear velocity. We use $K_{p}$ and $K_{d}$ gains of $20$ and $1$, respectively, which are used in the position controller. We again use action interpolation, an action range of 35% physical limits, and a $1$ step action history. We also use a second-order Butterworth low-pass filter with a high-cut value of $8$ to smooth the position targets. Finally, to reset the robot, we use the reset policy provided by (Smith et al., 2022). We train $4$ prior behavior policies for the real-world experiments, each of which is trained with a frozen knee joint. Specifically, we train a policy with the front right knee joint frozen, the front left knee joint frozen, the rear right knee joint frozen, and the rear left knee joint frozen. $\beta=0.5$ is used in all real-world experiments for ROAM.
@@
For RMA training, we changed the environment dynamics between each episode and trained for a total of $2,000,000$ iterations. The standard architecture and hyperparameter choices from (Kumar et al., 2021) were used along with DroQ (Hiraoka et al., 2021) as the base algorithm.